OpenStudy (anonymous):

find the derivative of 2x arccos(x)-2square root 1-x^2, i know that the answer is 2arccos x, help please

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

i see formula for the derivative, and i am not sure what to plug or combine like terms

OpenStudy (anonymous):

Let @Kainui help, Kainui ha??

OpenStudy (anonymous):

would you use productive rule?

OpenStudy (kainui):

I'm just watching good luck @OOOPS !

OpenStudy (anonymous):

@MrGrimm show me your work, please

OpenStudy (anonymous):

Of course we have to use product rule.

OpenStudy (dumbcow):

this may be helpful http://en.wikipedia.org/wiki/Inverse_trigonometric_functions#Derivatives_of_inverse_trigonometric_functions and yes you will be using the product rule as well as the chain rule

OpenStudy (dumbcow):

lets look at first term, using product rule \[\rightarrow (2x)' \cos^{-1} x + 2x(\cos^{-1} x)'\] \[= 2 \cos^{-1} x - \frac{2x}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

\[(-2\sqrt{1-x ^{2}})(2\arccos x)+(2xarccos x)(2\sqrt{1-x ^{2}})\]

OpenStudy (anonymous):

thats what i thought it would be

OpenStudy (dumbcow):

ok for 2nd term, we use chain rule \[\rightarrow \frac{-2 (-2x)}{2\sqrt{1-x^2}}\] because derivative of "1-x^2" is -2x simplifying to \[2 \cos^{-1} x -\frac{2x}{\sqrt{1-x^2}}+\frac{2x}{\sqrt{1-x^2}}\] \[= 2 \cos^{-1} x\]

OpenStudy (anonymous):

ok thats more clear thank you, was confused on what to take the derivative

OpenStudy (dumbcow):

yw